The ammonia molecule NH3 has a permanent electric dipole moment equal to 1.52 D, where 1 D = 1 debye unit = 3.34x10-30 C·m. Calculate the electric potential due to an ammonia molecule at a point 50.0 nm away along the axis of the dipole. Use V = 0 at infinity.  

Please if you do not know how to solve it just agnor it. I posted this question week ago and it was answered incorrectly. Therefore, my time and money were wasted
Answer #1

We have to find V in R. We'll suppose d<<r.

p=moment of the dipole

\vec{p}=q\vec{d}

q=charge

d=distance between the charges forming the dipole

The potential created by the dipole at distance r from the charges (considering the general case, in P; CP is another r):

V=\frac{kpcos\theta }{r^{2}}\: \left ( 1 \right )

k=\frac{1}{4\pi \varepsilon _{0}}

k=9*109 Nm2/C2

r=50*10-9 m

p~5.08*10-30 C*m

In the case of point R, alpha=0 --> cos=1.

Replace the values in eq(1) and get V~18.3*10-6V.